OpenStudy (anonymous):

ali has 6 cards the mea of the 6 cards is 10. the range of the 6 cards is 4 . what are the numbers on the other 2 cards cards:10 10 10 10 ? ?

OpenStudy (anonymous):

8 and 12

OpenStudy (amistre64):

thats better than my answer; i dint even understand the question :)

OpenStudy (anonymous):

thanks

OpenStudy (amistre64):

i was still trying to figure what cards:10 10 10 10 ?? meant lol

OpenStudy (anonymous):

If the mean of 6 cards is 10, then the sum of cards is 60. The sum of 4 cards are 4*10=40. So the other two cards make up 20. the range is 4, so, the maximum and min value differs by 4, x+y =20 and x-y=4. Solve for x and y

OpenStudy (anonymous):

see the range is 4 and total is 60 out of this the four cards of 10 total to 40, so 20 is left so we cant hv the two cards more than 10 or less than 10 as they will not total to 20 8, 10, 10, 10, 10, 12 give a range of 4 (12-8) and also 8+12 = 20

OpenStudy (anonymous):

will somebody give me a medal just found this site

OpenStudy (anonymous):

u get a medal for helping/answering and not for asking !!!!